# GS4: Generalizable Sparse Splatting Semantic SLAM

### v1

@@ 3.3 The SLAM System
An overview of the system is summarized in Fig. 2. The Gaussian prediction network predicts 3D Gaussians from RGB-D images, which can be rendered into high-fidelity color, depth, and semantic images. The Gaussian refinement network refines both the existing Gaussians in the 3D map and the newly added ones to accurately render both current and previous views. During testing, we occasionally run 1-iteration test-time optimization and refine 3D Gaussians in the map to reflect camera pose updates from loop closure and bundle adjustment in the tracking module.
@@ 3.3.1 Tracking and Global Bundle Adjustment
For camera tacking in our SLAM system, we adopt a tracker used in GO-SLAM [42] which is an enhanced version of DROID-SLAM’s tracking module [36]. In the Local Camera Tracking, a keyframe is initialized when sufficient motion is detected, and loop closure (LC) is performed. Meanwhile, the Global Localization is equipped with full bundle adjustment (BA) for real-time global refinement. Both LC and BA help address the problem of accumulated errors and drift that can occur during the localization process.
Local Camera Tracking. Local camera tracking processes RGB-D images to compute optical flow between each new frame and the previous keyframe [36]. If the average flow exceeds a predefined threshold, the current frame is selected as a new keyframe and added to the keyframe buffer for subsequent refinement. In the keyframe buffer, a loop candidate is accepted only if a frame pair mean flow is below a predefined threshold and the frames are far apart enough. Subsequently, a differentiable Dense Bundle Adjustment (DBA) [36] layer is applied to refine the camera pose for each keyframe.
Global Localization. Full bundle adjustment (BA) is performed in a separate thread, following strategies employed in [24]. This approach allows the system to continue tracking new frames and handling loop closures concurrently.
3.3.2 One-Iteration Gaussian Optimization
Loop closure and bundle adjustment are essential components in SLAM systems, employed to correct accumulated drift and adjust the camera poses of previous frames. However, these adjustments can cause Gaussians inserted based on earlier, uncorrected poses to misalign with the scene, leading to inaccurate rendering and mapping. It is crucial to implement a mechanism that updates the Gaussians in the 3D map following pose corrections. To address this issue, we propose using rendering-based optimization to update the Gaussian parameters , , and with only a single iteration. We render RGB-D images for the top-k frames, selected based on significant pose changes. This approach maintains the consistency of the 3D map with updated camera poses. To enhance the efficiency of this optimization, we employ the batch rendering technique from [39]. We omit semantic image rendering to improve system efficiency. For 1-iteration optimization, we add a SSIM term to the RGB loss, following [15]:
@@ 4.1 Experimental Setup
Training Settings. We train our Gaussian prediction and refinement networks entirely on RGB-D videos from the real ScanNet dataset. Excluding the six test scenarios commonly used in SLAM benchmarks, we utilize all remaining train and val scenarios for training, employing 20 common semantic classes for supervision. We use AutoFocusFormer [46] as the backbone for ODIN and for Gaussian prediction, we select the second stage of the backbone as the prediction stage. To train the Gaussian prediction network, we adopt a low-to-high resolution curriculum as described in [47], progressing through three stages with image resolutions of 256×256, 480×480, and 640×480. In the first two stages, training images are resized so that the shorter side is 256 and 480 pixels, respectively, and then center-cropped to a square. For the Gaussian refinement network, which operates across multiple consecutive frames, we use a multi-stage training strategy: starting with two consecutive frames, progressing to four, and finally using eight.
Datasets. During testing, we evaluate our method on six real-world scenes on ScanNet [7]. Additionally, we perform zero-shot experiments on several real scenes from NYUv2 [26] and TUM RGB-D [34], as well as synthetic scenes from the Replica dataset [32].
@@
Metrics. We use PSNR, Depth-L1 [44], SSIM [37], and LPIPS [41] to evaluate the reconstruction and rendering quality. We additionally report reconstruction metrics such as Accuracy, Completion, Completion Ratio (7cm) and F1 (7cm) in the supplemetary. For GS-based SLAM methods, we also show the number of Gaussians for each scene. For semantic segmentation, we report the mean Intersection over Union (mIoU). To evaluate the accuracy of the camera pose, we adopt the average absolute trajectory error (ATE RMSE) [33].
@@
Baselines. We compare our method to several state-of-the-art approaches, including NeRF-based SLAM methods such as NICE-SLAM [44], GO-SLAM [42] and Point-SLAM [31], as well as 3D Gaussian-based SLAM methods like SplaTAM [13], RTG-SLAM [28], GS-ICP SLAM [9], and Semantic 3D Gaussian-based SLAM such as SGS SLAM [18] and OVO-Gaussian-SLAM [21]. SGS-SLAM and OVO-Gaussian-SLAM are the only semantic SLAM methods available for comparison since the code is not available for other semantic SLAM approaches. Note that SGS-SLAM [18] and DNS-SLAM [17] employ test-time optimization using ground truth semantic labels on the testing set. And SGS-SLAM has been shown to outperform all other existing semantic SLAM methods [43, 17]. To ensure a fair comparison and simulating SLAM applications in real-world scenarios where ground truth semantic labels are unavailable, we trained a 2D segmentation model using a Swin backbone [20] with Mask2Former [5] on ScanNet, following the same training strategy as our model, and used predicted semantic labels to supervise SGS-SLAM.
@@ 4.2 Results
| Methods | Metrics | 0000 | 0059 | 0106 | 0169 | 0181 | 0207 | Avg |
| PSNR | 18.71 | 16.55 | 17.29 | 18.75 | 15.56 | 18.38 | 17.54 | |
| NICE-SLAM | SSIM | 0.641 | 0.605 | 0.646 | 0.629 | 0.562 | 0.646 | 0.621 |
| LPIPS | 0.561 | 0.534 | 0.510 | 0.534 | 0.602 | 0.552 | 0.548 | |
[…]
| Point-SLAM | SSIM | 0.806 | 0.765 | 0.676 | 0.686 | 0.823 | 0.750 | 0.751 |
| LPIPS | 0.485 | 0.499 | 0.544 | 0.542 | 0.471 | 0.544 | 0.514 | |
[…]
| SSIM | 0.589 | 0.792 | 0.748 | 0.776 | 0.615 | 0.694 | 0.702 | |
| SplaTAM | LPIPS | 0.485 | 0.289 | 0.314 | 0.281 | 0.456 | 0.358 | 0.364 |
| Depth L1 | 8.01 | 5.33 | 7.72 | 6.18 | 8.33 | 7.67 | 7.21 | |
| GS Num | 3089k | 2117k | 1821k | 1656k | 2714k | 3402k | 2466k | |
[…]
| SSIM | 0.375 | 0.291 | 0.401 | 0.408 | 0.390 | 0.367 | 0.372 | |
| RTG SLAM | LPIPS ↓ | 0.788 | 0.780 | 0.716 | 0.711 | 0.786 | 0.787 | 0.761 |
| Depth L1 | 90.5 | 116.00 | 77.40 | 92.50 | 91.90 | 117.03 | 97.56 | |
| GS Num | 1940k | 1201k | 1000k | 1101k | 1113k | 1021k | 1229k | |
[…]
| SSIM | 0.594 | 0.622 | 0.655 | 0.642 | 0.722 | 0.633 | 0.645 | |
| GS-ICP SLAM | LPIPS ↓ | 0.728 | 0.703 | 0.655 | 0.698 | 0.696 | 0.624 | 0.684 |
| Depth L1 | 218k | 259k | 236k | 223k | 216k | 196k | 224k | |
| GS Num | 5666k | 2586k | 2506k | 2190k | 2519k | 3027k | 2565k | |
[…]
| SSIM | 0.511 | 0.742 | 0.636 | 0.572 | 0.531 | 0.569 | 0.594 | |
| SGS SLAM | LPIPS | 0.687 | 0.485 | 0.597 | 0.613 | 0.706 | 0.600 | 0.615 |
| Depth L1 | 9.11 | 6.73 | 10.66 | 18.89 | 15.02 | 10.56 | 11.83 | |
| GS Num | 2840k | 2020k | 1582k | 1266k | 2054k | 2926k | 2114k | |
[…]
| SSIM | 0.864 | 0.858 | 0.794 | 0.871 | 0.859 | 0.853 | 0.850 | |
| GS4 (Ours) | LPIPS | 0.348 | 0.311 | 0.366 | 0.284 | 0.377 | 0.375 | 0.344 |
| Depth L1 | 3.89 | 6.06 | 9.85 | 7.66 | 4.30 | 4.70 | 6.08 | |
| GS Num | 438k | 227k | 220k | 241k | 246k | 397k | 295k |
| Methods | Metrics | 0000 | 0059 | 0106 | 0169 | 0181 | 0207 | Avg |
| PSNR | 20.76 | 16.24 | 18.11 | 14.93 | 17.99 | 21.20 | 18.21 | |
| SSIM | 0.784 | 0.606 | 0.685 | 0.407 | 0.694 | 0.766 | 0.657 | |
| GO-SLAM | LPIPS | 0.497 | 0.562 | 0.556 | 0.600 | 0.601 | 0.502 | 0.553 |
| Depth L1 | 6.58 | 21.38 | 15.87 | 48.27 | 8.89 | 7.85 | 18.14 | |
[…]
| SSIM | 0.901 | 0.875 | 0.808 | 0.884 | 0.881 | 0.884 | 0.872 | |
| GS4 (Ours) | LPIPS | 0.238 | 0.257 | 0.307 | 0.257 | 0.284 | 0.268 | 0.269 |
| Depth L1 | 3.93 | 6.16 | 9.68 | 7.09 | 4.40 | 4.97 | 6.04 | |
| GS Num | 194k | 99k | 110k | 112k | 127k | 177k | 136k |
@@
Rendering and Reconstruction Performance. In Table 4.2, we evaluate the rendering and reconstruction performance of our SLAM method on ScanNet. This is a difficult task compared to the synthetic data where neural RGB-D SLAM methods usually show strong results, because inevitably inaccurate ground truth camera poses and depths make optimization much harder than completely clean synthetic datasets. Compared to existing dense neural RGB-D SLAM methods, our method achieves the best performance in PSNR, SSIM, and Depth L1 metrics. Specifically, our method surpasses the runner-up, Point-SLAM [30], by 2.89 dB in PSNR (a 14.6 percent improvement) and 0.091 in SSIM (a 13.2 percent improvement). Furthermore, our approach utilizes approximately 10x fewer Gaussians than the baselines. This efficiency highlights the effectiveness of our method in achieving high-quality scene representation with reduced computational complexity.
In Table 4.2, we ran our method with input to compare against GO-SLAM which shares the same tracking method as ours but renders at the same low resolution. GS4 maintains the same PSNR and depth prediction quality as its high resolution version and significantly outperforms GO-SLAM across all metrics.
Fig. 3 shows visual results of RGB and depth rendering. Our method demonstrates superior performance than other GS-SLAM methods. Notably, sometimes the depth maps of our approach even outperforms the noisy ground truth depth inputs. For instance, in the first two columns in Fig. 3, our method delivers a more contiguous and complete rendering of the bicycle tires. Similarly, in the middle two columns in Fig. 3, we reconstruct the chair’s backrest nearly entirely, whereas the GT depth data lacks this detail.
| Methods | SGS SLAM | OVO-Gaussian-SLAM | GS4 (Ours) |
| mIoU(2D) | 37.20 | ✗ | 60.40 |
| mIoU(3D) | 18.87 | 32.58 | 50.00 |
@@
Semantic Performance. In Table 3, we present both 2D rendering and 3D mean Intersection over Union (mIoU) scores across the six ScanNet test scenes. For 3D mIoU evaluation, we first align the reconstructed map with the ground-truth mesh and then use 3D neighborhood voting to assign predicted labels. Our method outperforms the previous runner-up, OVO-Gaussian-SLAM, by 17.42 in 3D mIoU. In terms of 2D rendering mIoU, our method surpasses the state-of-the-art semantic GS-based SLAM approach, SGS-SLAM, by 23.2. As illustrated in Fig. 5, our approach achieves superior semantic segmentation accuracy compared to the SGS-SLAM baseline. For example , in the first column of Fig.5, our semantic rendering provides a more accurate representation of the desks, chairs, and night tables than SGS-SLAM.

### v2

@@ 3.3 The SLAM System
An overview of the system is summarized in Fig. 3. The system always maintains a set of 3D Gaussians representing the entire scene. For each new RGB-D image, the Gaussian prediction network predicts 3D Gaussian parameters, which can be rendered into high-fidelity color, depth, and semantic images. The Gaussian refinement network refines both the existing Gaussians in the 3D map and the newly added ones to accurately render both current and previous views. During testing, we occasionally run few-iteration test-time optimization and refine 3D Gaussians in the map to reflect camera pose updates from loop closure and bundle adjustment in the tracking module.
@@ 3.3.1 Tracking and Global Bundle Adjustment
For camera tacking in our SLAM system, we adopt a tracker used in GO-SLAM (Zhang et al., 2023) which is an enhanced version of DROID-SLAM’s tracking module (Teed & Deng, 2021). It first predicts motion in every frame. In local camera tracking, a keyframe is initialized when sufficient motion is detected, and loop closure (LC) is performed. Meanwhile, global localization performs full bundle adjustment (BA) for real-time global refinement once the system contains more than 25 keyframes. Both LC and BA help address the problem of accumulated errors and drift that can occur during the localization process.
3.3.2 Few-Iteration Gaussian Optimization
Loop closure and bundle adjustment are essential components in SLAM systems, employed to correct accumulated drift and adjust the camera poses of previous frames. However, these adjustments can cause Gaussians inserted based on earlier, uncorrected poses to misalign with the scene, leading to inaccurate rendering and mapping. It is crucial to implement a mechanism that updates the Gaussians in the 3D map following pose corrections. To address this issue, we propose using rendering-based optimization to update the Gaussian parameters , , and with only a few iterations. We render RGB-D images for the top-k frames, selected based on significant pose changes. This approach maintains the consistency of the 3D map with updated camera poses. To enhance the efficiency of this optimization, we employ the batch rendering technique from Ye et al. (2024). We omit semantic image rendering to improve system efficiency. For few-iteration optimization, we add a SSIM term to the RGB loss, following Kerbl et al. (2023b):
@@ 4.1 Experimental Setup
Training Settings. We train our Gaussian prediction and refinement networks entirely on RGB-D videos from the real ScanNet datasets We exclude the six standard SLAM test scenarios and use all remaining training and validation scenes, supervising with 20 common semantic classes. For ScanNet++v2, we exclude the two SLAM test scenarios and use all other available training scenes, supervising with the 100 most common semantic classes. We adopt AutoFocusFormer (Ziwen et al., 2023) as the backbone for both Mask2Former and Gaussian prediction, using the second stage of the backbone as the prediction stage. Additionally, we experiment with Swin Transformer (Liu et al., 2021) as an alternative backbone. The detailed results are provided in the appendix. Following the low-to-high resolution curriculum of (Ziwen et al., 2024), we train the Gaussian prediction network in three stages with input resolutions of 256×256, 480×480, and 640×480. In the first two stages, images are resized such that the shorter side is 256 or 480 pixels and then center-cropped to a square. On ScanNet++v2, we use resolutions of 256x384. For the refinement network, which processes multiple consecutive frames, we adopt a progressive training schedule: beginning with two frames, then four, and finally eight.
Datasets. During testing, we evaluate our method on six real-world scenes on ScanNet (Dai et al., 2017a) , and six real-world scenes on ScanNet++ (Yeshwanth et al., 2023). ScanNet++ has 11 scenes with continuous camera trajectory, but SplaTAM and SGS-SLAM completely fail in 5 scenes, respectively. Hence we only use the 6 scenes they can finish running, so that their numbers are not infinitely bad. Additionally, we perform zero-shot experiments on real scenes from NYUv2 (Nathan Silberman & Fergus, 2012) and TUM RGB-D (Sturm et al., 2012a).
@@
Metrics. We use PSNR, Depth-L1 (Zhu et al., 2022a), SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) to evaluate the reconstruction and rendering quality. We additionally report reconstruction metrics such as Accuracy, Completion, Completion Ratio (7cm) and F1 (7cm) in the appendix. For GS-based SLAM methods, we also report the number of Gaussians. For semantic segmentation, we report the mean Intersection over Union (mIoU). To evaluate the accuracy of the camera pose, we adopt the average absolute trajectory error (ATE RMSE) (Sturm et al., 2012b).
@@
Baselines. We compare our method against several state-of-the-art approaches: NeRF-based SLAM methods, including NICE-SLAM (Zhu et al., 2022a), GO-SLAM (Zhang et al., 2023), and Point-SLAM (Sandström et al., 2023); 3D Gaussian-based SLAM methods such as SplaTAM (Keetha et al., 2024), RTG-SLAM (Peng et al., 2024), and GS-ICP SLAM (Ha et al., 2024); and semantic 3D Gaussian-based SLAM methods, including SGS-SLAM (Li et al., 2024b), GS3LAM (Li et al., 2024a) and OVO-Gaussian-SLAM (Martins et al., 2024). SGS-SLAM, GS3LAM and OVO-Gaussian-SLAM are the only semantic SLAM methods available for comparison since the code is not available for other semantic SLAM approaches. Note that SGS-SLAM (Li et al., 2024b) and GS3LAM (Li et al., 2024a) employ test-time optimization using ground truth semantic labels on the test set. SGS-SLAM and GS3LAM have been shown to outperform all other existing semantic SLAM methods (Zhu et al., 2024; Li et al., 2023). To ensure a fair comparison and simulating SLAM applications in real-world scenarios where ground truth semantic labels are unavailable, we trained a 2D segmentation model using a Swin backbone (Liu et al., 2021) with Mask2Former (Cheng et al., 2022) on ScanNet and ScanNet++, following the same training strategy as our model, and used predicted semantic labels to supervise SGS-SLAM nad GS3LAM.
@@ 4.2 Results
Rendering and Reconstruction Performance. In Table 1, we evaluate the rendering and reconstruction performance of our method on ScanNet. This is a difficult task compared to the synthetic data where neural RGB-D SLAM methods usually show strong results, because inevitably inaccurate ground truth camera poses and depths make optimization much harder than completely clean synthetic datasets. Compared to existing dense neural RGB-D SLAM methods, our approach achieves state-of-the-art performance on PSNR, SSIM, and Depth L1 metrics. Specifically, our method surpasses the runner-up, GS3LAM (Li et al., 2023), by 3.59 dB in PSNR (a 17.4 percent improvement). and 0.091 in SSIM (a 10.4 percent improvement). Furthermore, our approach utilizes approximately 10x fewer Gaussians than the baselines. This efficiency highlights the effectiveness of our method in achieving high-quality scene representation with reduced computational complexity.
In Table 1, we ran our method with input to compare against GO-SLAM which shares the same tracking method as ours but renders at the same low resolution. GS4 maintains the same PSNR and depth prediction quality as its high resolution version and significantly outperforms GO-SLAM across all metrics.
Fig. 4 shows visual results of RGB and depth rendering. Our method demonstrates superior performance than other GS-SLAM methods. Notably, sometimes the depth maps of our approach even turn out to be better than the noisy ground truth depth inputs. For instance, in the first two columns, our method delivers a more contiguous and complete rendering of the bicycle tires. Similarly, in the middle two columns, we reconstruct the chair’s backrest nearly entirely, whereas the GT depth data lacks this detail.
| Res | Method | PSNR | SSIM | LPIPS | Depth L1 | GS Num |
| NICE-SLAM | 17.54 | 0.621 | 0.548 | - | - | |
| Point-SLAM | 19.82 | 0.751 | 0.514 | - | - | |
| SplaTAM | 18.99 | 0.702 | 0.364 | 7.21 | 2466k | |
| RTG SLAM | 12.75 | 0.372 | 0.761 | 97.56 | 1229k | |
| GS-ICP SLAM | 14.73 | 0.645 | 0.684 | 103.31 | 2565k | |
| SGS SLAM | 15.89 | 0.594 | 0.615 | 11.83 | 2114k | |
| GS3LAM | 20.67 | 0.796 | 0.288 | 11.88 | 2154k | |
| GS4 (Ours, AFF, 1 iter) | 22.61 | 0.850 | 0.335 | 6.55 | 356k | |
| GS4 (Ours, AFF, 5 iters) | 24.26 | 0.879 | 0.304 | 4.98 | 245k | |
| GO-SLAM | 18.21 | 0.657 | 0.553 | 18.14 | - | |
| GS4 (Ours, AFF, 1 iter) | 22.50 | 0.885 | 0.238 | 6.04 | 172k | |
| GS4 (Ours, AFF, 5 iters) | 24.02 | 0.915 | 0.201 | 5.05 | 130k |
@@
Semantic Performance. In Table 2, we present both 2D rendering and 3D mean Intersection over Union (mIoU) scores across the six ScanNet test scenes. For 3D mIoU evaluation, we first align the reconstructed map with the ground-truth mesh and then use 3D neighborhood voting to assign predicted labels. Our method outperforms the previous runner-up, GS3LAM, by 20.42 in 3D mIoU and by 6.62 in 2D mIoU. Qualitative comparisons are provided in the appendix.
| Methods | SGS SLAM | OVO-Gaussian-SLAM | GS3LAM | GS4 (Ours, AFF, 1 iter) | GS4 (Ours, AFF, 5 iters) |
| mIoU(2D) | 37.20 | ✗ | 56.42 | 63.04 | 62.36 |
| mIoU(3D) | 18.87 | 32.58 | 34.42 | 54.84 | 53.61 |
Tracking Performance. Table 3 shows the tracking results. Our method uses the same tracking algorithm as GO-SLAM, which is significantly better than other GS-based SLAM methods.
| Metric | NICE-SLAM | Point-SLAM | SplaTAM | RTG SLAM | GS-ICP SLAM | SGS SLAM | GS3LAM | GO-SLAM | GS4 (Ours) |
| ATE RMSE [cm] | 10.70 | 12.19 | 11.88 | 144.52 | NaN | 40.97 | 30.88 | 7.00 | 6.98 |
Runtime Comparison. Table 4 presents a runtime comparison of our method against the baselines at the resolution, conducted on an Nvidia RTX TITAN. FPS is calculated by dividing the total number of frames by the total time to represent the overall system performance. While GS-ICP SLAM is faster than ours, its rendering and tracking performance is significantly worse (Table 1 and 3). Our approach is 12x faster than SplaTAM, 17x faster than SGS-SLAM, and 24x faster than GS3LAM.
| Methods | Point-SLAM | SplaTAM | RTG-SLAM | GS-ICP SLAM | SGS-SLAM | GS3LAM | GS4 (ours, AFF, 1 iter) | GS4 (ours, AFF, 5 iters) |
| FPS | 0.05 | 0.23 | 1.01 | 3.62 | 0.17 | 0.12 | 2.87 | 1.92 |
ScanNet++ Experiments. Table 5 presents the 2D rendering results for both training and novel views, as well as the 3D semantic, tracking, and runtime comparisons of our method against the baselines at a resolution of 256384. The results are averaged over six scenes. Our method achieves state-of-the-art performance in both 2D and 3D tasks, showing a 10 improvement in PSNR and SSIM on the training view and a 13.22 improvement in 3D mIoU over the runner-up, while using only 2.0–2.5 of the number of Gaussians required by other methods. Compared to SGS-SLAM, although it achieves competitive 2D mIoU on the training view, its performance drops drastically on the novel view (2.46) and in 3D mIoU (0.58), indicating severe overfitting to the input views and poor generalization to the full 3D scene. In addition to delivering superior performance across both 2D and 3D metrics, our method also exhibits substantially higher efficiency, running 19× faster than SplaTAM and 25× faster than SGS-SLAM.
| Method | Train View | Novel View | mIoU (3D) | GS Num | ATE RMSE | FPS | ||||||||
| PSNR | SSIM | LPIPS | Depth L1 | mIoU (2D) | PSNR | SSIM | LPIPS | Depth L1 | mIoU (2D) | |||||
| SplaTAM | 22.98 | 0.854 | 0.218 | 7.11 | ✗ | 17.70 | 0.662 | 0.444 | 10.28 | ✗ | ✗ | 1899k | 1588.95 | 0.17 |
| SGS-SLAM | 21.29 | 0.801 | 0.276 | 10.52 | 28.24 | 9.70 | 0.265 | 0.664 | 86.93 | 2.46 | 0.58 | 2374k | 1315.89 | 0.13 |
| GS4 (Ours) | 25.26 | 0.929 | 0.162 | 21.93 | 23.51 | 15.61 | 0.531 | 0.425 | 31.90 | 13.27 | 13.80 | 48k | 5.88 | 3.23 |
Zero-shot Experiments. In Table 6, we report quantitative zero-shot results. For NYUv2, the numbers are averaged over three scenes, and for TUM-RGBD, they are also averaged over three scenes. Per-scene results are provided in the appendix. On NYUv2, our method outperforms all other GS-based SLAM approaches across all rendering metrics while using significantly fewer Gaussians. Qualitative comparisons are also provided in the appendix. On TUM-RGBD, our method outperforms the baselines in terms of SSIM and the number of Gaussians, and closely matches the best performance in other metrics, despite relying primarily on a feed-forward model trained on ScanNet.
| Dataset | Res | Method | PSNR | SSIM | LPIPS | GS Num |
| SplaTAM | 18.86 | 0.692 | 0.372 | 1236k | ||
| NYUv2 | RTG-SLAM | 11.84 | 0.221 | 0.703 | 807k | |
| SGS-SLAM | 19.32 | 0.708 | 0.357 | 1108k | ||
| GS4 (Ours, AFF, 5 iters) | 22.09 | 0.853 | 0.268 | 278k | ||
| SplaTAM | 22.76 | 0.891 | 0.182 | 803k | ||
| TUM RGBD | RTG-SLAM | 19.75 | 0.769 | 0.395 | 198k | |
| SGS-SLAM | 22.44 | 0.876 | 0.184 | 735k | ||
| GS4 (Ours, AFF, 5 iters) | 22.70 | 0.903 | 0.191 | 166k |
Ablation Study. We conduct an ablation study using all ScanNet test scenes, as shown in Table 7. The results demonstrate that both the Gaussian Refinement Network and the Few-Iteration Gaussian Optimization are critical to the performance of GS4. Additionally, Gaussian pruning significantly reduces the number of Gaussians without sacrificing accuracy. More results on backbone comparisons, reconstruction accuracy and qualitative results are shown in the appendix.
| Design Choice | PSNR [dB] | SSIM | LPIPS | Depth L1 | mIoU | Gs Num |
| GS Prediction | 14.91 | 0.460 | 0.663 | 33.14 | 41.0 | 133k |
| + GS Refinement | 16.15 | 0.556 | 0.584 | 29.56 | 44.68 | 576k |
| + Few-Iter. Optimization (1) | 22.66 | 0.852 | 0.335 | 6.41 | 63.7 | 680k |
| + GS Pruning (Full SLAM) | 22.61 | 0.851 | 0.335 | 6.56 | 63.1 | 355k |
5 Conclusion
We present GS4, a novel SLAM system that incrementally constructs and updates a 3D semantic scene representation from a monocular RGB-D video with a learned generalizable network. Our novel Gaussian refinement network and few-iteration Gaussian optimization significantly improve the performance of our approach. Our experiments demonstrate state-of-the-art semantic SLAM performance on the ScanNet benchmark while running 10x faster and using 10x less Gaussians than baselines. The model also showed strong generalization capabilities through zero-shot transfer to the NYUv2 and TUM RGB-D datasets. In future work, we will further improve the computational speed of GS4 and explore options for a pure RGB-based SLAM approach.
Appendix
Appendix A More Experimental Setup
For ScanNet, we evaluate on six scenes (scene0000, scene0059, scene0106, scene0169, scene0181, and scene0207), which are commonly used by other SLAM methods. For ScanNet++, we use six scenes in total: 8b5caf3398, 3f15a9266d, e7af285f7d, 99fa5c25e1, 09c1414f1b, and 9071e139d9.
Appendix B Reconstruction Results on ScanNet
In Table 9, we use the metrics including Accuracy (Acc.), Completion (Comp.), Completion Ratio[7cm] and F-Score[7cm] to evaluate the scene geometry on ScanNet. The definitions of the evaluation metrics are detailed in Table 8. GS4 outperforms all baselines in terms of completion and F-score.
| 3D Metric | Formula |
| Acc | |
| Comp | |
| Completion Ratio[7cm] | |
| F-score[7cm] |
| Methods | Acc. | Comp. | Comp. Ratio (<7cm) | F-Score (<7cm) | GS Num |
| SplaTAM | 8.10 | 5.58 | 76.34 | 75.95 | 2466k |
| RTG-SLAM | 99.80 | 47.44 | 24.61 | 16.69 | 1229k |
| SGS-SLAM | 17.11 | 13.75 | 55.01 | 55.26 | 2114k |
| GS4 | 8.56 | 3.88 | 87.34 | 79.90 | 295k |
Appendix C Qualitative Comparison for Semantic Segmentation.
As illustrated in Fig. 5, our approach achieves superior semantic segmentation accuracy compared to the SGS-SLAM baseline. For example, in the first column of Fig.5, our semantic rendering provides a more accurate representation of the desks, chairs, and night tables than SGS-SLAM.
SGS SLAM |
GS4 (Ours) |
GT (labels) |
GT (RGB) |
| scene0000 | scene0169 | scene0207 |
Appendix D Zero-Shot Results on NYUv2
Fig. 6 illustrates our zero-shot visualization results on the NYUv2 dataset. Despite our models being exclusively trained on the ScanNet dataset, our method demonstrates superior performance on the NYUv2 dataset compared to other GS-based SLAM approaches. In Table 10, we present the quantitative zero-shot results across three scenes from the NYUv2 dataset. Our method outperforms all other GS-based SLAM approaches on all rendering metrics, while using significantly fewer Gaussians.